# Invariance principle in dynamical systems

### v1

@@ 1. Introduction
We can relate foliations and invariant measures via disintegration, see Section 2 for the definitions and properties of the disintegration. Understanding these disintegrations gives a powerful tool to study ergodic properties of our dynamical systems. An example of this are the SRB measures of hyperbolic systems defined in [Sin72], [Rue76], [Bow08]. These measures capture the statistical behavior of almost all points, with respect to the Lebesgue measure and they are characterized by its disintegration along the unstable foliation.
@@
In this survey we will focus on partially hyperbolic dynamical systems that admits a center invariant foliation and understand what is the relation between entropy and the disintegration along this foliation. The result we are going to explain shows that if the center foliation “does not carry entropy” then the disintegration of the measure is invariant under the holonomies of the stable and unstable foliations. This is known as the Invariance Principle.
@@
We denote ${\mathcal{Q}}(x)$ as the element of $M$ containing the point $x$. If we have two partitions $\mathcal{Q}$} and $\mathcal{Q}$} such that ${\mathcal{Q}}(x)\subset{\mathcal{P}}(x)$ for all $x\in M$, we can write ${\mathcal{Q}}\geq{\mathcal{P}}$.
@@
A $C^{1}$ diffeomorphism $f\colon M\to M$ is partially hyperbolic if there exist a nontrivial splitting of the tangent bundle
@@
The main idea is to use the definition of unstable entropy and Jensen inequality for convex functions to find an inequality between unstable entropy and the full entropy, then use that the equality in Jensen is only satisfy when the function is constant almost everywhere, this will give a condition on the disintegration that will be translated to $u$-invariance.
@@ 5. Random dynamics and Furstenberg entropy
By Oseledets theorem [Ose68], the limit is well defined for $\nu^{\mathbb{Z}}\times\eta$-almost every $(\omega,x)$. Let $\lambda_{min}(\omega,x)$ be the smallest Lyapunov exponent on $(\omega,x)$.

### v2

@@ 1. Introduction
We can relate foliations and invariant measures via disintegration, see Section 2 for the definitions and properties of the disintegration. Understanding these disintegrations gives a powerful tool to study ergodic properties of our dynamical systems. An example of this are the SRB measures of hyperbolic systems defined in [Sin72], [Rue76], [Bow08]. These measures capture the statistical behavior of almost all points, with respect to the Lebesgue measure and they are characterized by their disintegration along the unstable foliation.
@@
In this survey we will focus on partially hyperbolic dynamical systems that admit a center invariant foliation and we will understand what is the relation between entropy and the disintegration along this foliation. The result we are going to explain shows that if the center foliation “does not carry entropy” then the disintegration of the measure is invariant under the holonomies of the stable and unstable foliations. This is known as the Invariance Principle.
@@
Given two partitions $\mathcal{P}$} and $\mathcal{Q}$} such that ${\mathcal{Q}}(x)\subset{\mathcal{P}}(x)$ for all $x\in M$, we write ${\mathcal{Q}}\geq{\mathcal{P}}$.
@@
A $C^{1}$ diffeomorphism $f\colon M\to M$ is partially hyperbolic if there exists a nontrivial splitting of the tangent bundle
@@
The main idea is to use the definition of unstable entropy and Jensen inequality for convex functions to find an inequality between unstable entropy and the full entropy, then use that the equality in Jensen is only satisfied when the function is constant almost everywhere, this will give a condition on the disintegration that will be translated to $u$-invariance.
@@ 5. Random dynamics and Furstenberg entropy
By Oseledets theorem [Ose68], the limit is well defined for $\nu^{\mathbb{N}}\times\eta$-almost every $(\omega,x)$. Let $\lambda_{min}(\omega,x)$ be the smallest Lyapunov exponent on $(\omega,x)$.